Selected T-accounts of Moore Company are given below for the just completed year:

Raw Materials
Bal. 1/1 15,000 Credits ?
Debits 120,000
Bal. 12/31 25,000
Manufacturing Overhead
Debits 230,000 Credits ?
Work in Process
Bal. 1/1 20,000 Credits 470,000
Direct materials 90,000
Direct labor 150,000
Overhead 240,000
Bal. 12/31 ?
Factory Wages Payable
Debits 185,000 Bal. 1/1 9,000
Credits 180,000
Bal. 12/31 4,000
Finished Goods
Bal. 1/1 40,000 Credits ?
Debits ?
Bal. 12/31 60,000
Cost of Goods Sold
Debits ?

Required:

1. What was the cost of raw materials used in production during the year?

2. How much of the materials in (1) above consisted of indirect materials?

3. How much of the factory labor cost for the year consisted of indirect labor?

4. What was the cost of goods manufactured for the year?

5. What was the unadjusted cost of goods sold for the year? Do not include any underapplied or overapplied overhead in your answer.

6. If overhead is applied to production on the basis of direct labor cost, what predetermined overhead rate was in effect during the year?

7. Was manufacturing overhead underapplied or overapplied? By how much?

8. Compute the ending balance in Work in Process. Assume that this balance consists entirely of goods started during the year. If $8,000 of this balance is direct labor cost, how much of it is direct materials cost? Applied overhead cost?
Answer #1

1) Cost of raw material used = 15000+120000-25000 = 110000

2) Indirect material = 110000-90000 = 20000

3) Indirect labor = 180000-150000 = 30000

4) Cost of goods manufactured = 470000

5) Unajdusted cost of goods sold = 40000+470000-60000 = 450000

6) Overhead rate = 240000/150000 = 160% of labor cost

7) Applied overhead = 240000

Actual overhead = 230000

Over applied overhead = 240000-230000 = 10000

8) Ending work in process = 30000

Applied overhead = 8000*1.6 = 12800

Direct material = 30000-20800 = 9200